# Quantitative LLM Judges

### v1

@@ Abstract
LLM-as-a-judge is a framework in which a large language model (LLM) automatically evaluates the output of another LLM. We propose quantitative LLM judges, which align evaluation scores of existing LLM judges to human scores in a given domain using regression models. The models are trained to improve the score of the original judge by using the judge’s textual evaluation and score. We present four quantitative judges for different types of absolute and relative feedback, which showcases the generality and versatility of our framework. Our framework is more computationally efficient than supervised fine-tuning and can be more statistically efficient when human feedback is limited, which is expected in most applications of our work. We validate these claims empirically on four datasets using two base judges. Our experiments show that quantitative judges can effectively improve the predictive power of existing judges through post-hoc modeling.
@@ 1 Introduction
Measuring the quality of generated natural language presents several challenges due to the diverse range of generation methods and evaluation criteria [10]. The LLM-as-a-judge paradigm has recently emerged as a compelling approach to these evaluation challenges. By leveraging the reasoning capabilities of large language models (LLMs), this approach can provide more nuanced assessments that correlate better with human judgments across diverse tasks [13]. Many LLM judges generate both qualitative feedback and numerical scores, thus combining the comprehensiveness of human evaluation with the scalability of automated metrics.
@@
Recent studies highlight issues with LLM judges such as low alignment with human scores, miscalibration, score compression, high variance, prompt sensitivity, and leniency bias [39, 46]. To address these issues, various approaches have been proposed to fine-tune LLMs to improve score prediction on specific tasks [9, 28]. However, fine-tuning still requires gradient updates across billions of model parameters, and the associated data and compute costs often make this approach impractical. As a result, many users continue to rely on off-the-shelf models like GPT-4 or Claude [2, 15].
@@
A key source of the issues with current LLM judges is that they conflate qualitative reasoning with quantitative assessment. For example, fine-tuned LLM judges like Prometheus [21] are trained with the cross-entropy (CE) loss on gold-standard textual evaluations to generate both qualitative feedback and numerical scores. This is an inherent mismatch in objectives, as LLMs optimized for next-token prediction are tasked with producing accurate numerical scores, a fundamentally different statistical problem. While LLMs excel at producing structured textual evaluations and semantic embeddings, they are poorly suited for regressing human scores or preferences [9].
@@
This mismatch raises a natural question: can qualitative summarization be decoupled from quantitative score prediction to achieve more accurate judgment? By decoupling these objectives, the LLM can focus on its strength—producing textual evaluations through reasoning—while accurate numerical score prediction is handled by classic machine learning models that are robust and loved by practitioners. This perspective is supported by prior works in interpretability and probing, which show that when model representations encode information relevant for downstream tasks, simple linear decoders can recover this information effectively [1, 14, 16, 3].
@@
Building on this insight, we introduce quantitative judges, a framework that enhances the original base judge by using its textual evaluation to predict more accurate numerical scores. We formulate four different quantitative judges for absolute rating and relative preference prediction tasks. Each judge has two stages: in the qualitative stage, a frozen LLM judge generates a textual evaluation and initial score, and in the quantitative stage, these outputs are used to predict a better score. Our design is general, efficient, and applies to any base LLM judge. Specifically:
@@
    General: Our judges predict human scores using the base judge’s textual evaluation and score. The predictor is a generalized linear model (GLM) [47] trained on human scores in the domain. The judges can be applied to absolute rating prediction tasks, such as regression and classification, or relative preference prediction tasks. They can be also applied to any base judge because we treat it as a black box. We show the versatility of our framework by proposing four quantitative judges.
@@
    Statistically efficient: Our judges are based on GLMs, which can be learned from limited feedback. This is expected in most applications of our work. Our judges are designed such that the base judge’s score or a distribution of it serves as a bias term in our model. This allows us, at least in principle, to always learn at least as good judge as the base judge.
@@
    Computationally efficient: Learning of classic machine learning models on the top of LLM embeddings is more computationally efficient, and can also be more statistically efficient, than fine-tuning of LLMs. We report an order of magnitude speedups in Section 5.
We comprehensively evaluate all proposed quantitative judges, showing their strengths and weaknesses. We also conduct ablation studies on training set size, regularization strength, and embedding choices. Through experiments on two base judges. Prometheus [21] and Llama 3.1 [12], we demonstrate that our quantitative judges consistently outperform base judges on both absolute rating and relative preference datasets. They can also outperform fine-tuning of the base judges on both quality and computational efficiency simultaneously. In summary, quantitative judges offer a practical and effective alternative to fine-tuning for improving LLM-based evaluation.
@@ 2 Background
LLMs are increasingly used not only to generate outputs but also to evaluate them, a paradigm known as the LLM-as-a-judge [21, 41, 13, 53, 39, 56]. This approach uses natural language capabilities of LLMs to simulate human evaluations, offering a scalable, cost-effective, and reproducible alternative. The output of the LLM judge can be a natural language assessment, a quantitative score, or a pairwise preference. Both the text assessment and a numeric score are often generated [21]. While efficient, the metrics that LLM judges produce have been criticized as not consistently aligned with human assessments [8, 10].
2.1 LLM-as-a-Judge and Fine-Tuning
@@
Early methods, such as Prometheus [21], introduced LLMs fine-tuned for both absolute rating and pairwise ranking. These models tried to replicate human judgment, often outperforming heuristic metrics in aligning with human preferences. Recent works, like LLMEval [53] and others [43, 25], further refined this setup by training on diverse instruction-response datasets, such as Alpaca52k [4], LMSYS-Chat [54], and ToxicChat [27], using feedback from GPT-4 or GPT-3.5. To improve the consistency and depth of judging, CritiqueLLM [19] proposed multi-path prompting that combines pointwise and pairwise strategies and JudgeLM [56] explored reference-free fine-tuning.
@@
Self-curation methods have been proposed to improve the alignment of judges. Yuan et al. [50] introduced self-rewarding LMs that generate both responses and reward signals to create preference datasets for direct preference optimization (DPO) [34]. Pang et al. [31] and Trivedi et al. [41] explored rationalization-based preferences by generating multiple CoT explanations from a fixed seed model. Other works [49, 52] integrated iterative feedback from stronger models or humans to refine evaluation granularity. While these methods improve alignment, they often rely on aggressive fine-tuning, which can introduce variance and bias. Small changes in the training data can lead to noticeable shifts in outputs, particularly when human feedback is limited. Moreover, strong priors from pre-training may be inadvertently distorted, undermining judgment consistency.
2.2 LLM Benchmarks and Calibration Challenges
@@
Evaluating the effectiveness of LLM judges has become an active area of study. JudgeBench [38] and Eval4NLP [51] proposed benchmarks where LLMs evaluate paired responses, aiming to reflect model reasoning and preference ranking. BigGenBench [20] expanded this by covering $77$ diverse tasks, although with limited instances per task, affecting statistical robustness. Thakur et al. [39] showed that while large judges align well with humans, smaller models often under-perform. To reduce evaluation variance, some methods compute mean scores by averaging outputs across multiple runs or implement voting schemes. FLEUR [22] integrated logit-based probability weighting to enhance score reliability for multi-modal evaluation. However, evaluations still suffer from dataset limitations and inconsistent generalization, especially since judges are often trained on general world knowledge and not calibrated to the particular domain being evaluated.
2.3 Drawbacks of LLM-as-a-Judge
@@
Recent studies have surfaced critical limitations in LLM-as-a-judge systems. Goel et al. [11] found that the judges tend to favor models similar in architecture or training data, creating blind spots due to shared inductive biases. Li et al. [24] revealed the preference leakage problem, where evaluation outcomes are skewed by overlap between judge training data and model-generated responses. The issues such as overconfidence [40, 7, 48] and bias [30] persist even after extensive fine-tuning, which additionally decreases the model’s generalization [17]. Despite ongoing efforts using reinforcement learning from human feedback (RLHF), robust, unbiased, and well-calibrated evaluation remains elusive [23, 18, 26]. As noted in a recent survey [13], LLM evaluators are still unreliable in high-stakes or open-ended judgment scenarios.
@@ 3 Setting
We study two types of LLM judges: evaluating a single response and comparing two responses.
Absolute LLM judge: The absolute judge evaluates a single LLM response. The evaluation can have various forms: text, score, or both. The score can evaluate various aspects of the response, such as coherence, correctness, factual consistency, relevance, and adherence to task-specific guidelines. In this work, we assume that the judge generates both a textual evaluation and its score. Specifically, let $(x,y)$ be a prompt-response pair from a judged LLM. The absolute judge maps $(x,y)$ to $(e,b)$, where $e$ is the textual evaluation of $y$ given $x$ and $b\in\mathbb{R}$ is its score, reflecting its absolute value.
The primary advantage of an absolute judge is its consistency and standardization in scoring across different responses. However, it may require extensive prompt engineering or fine-tuning to align with human evaluations [18].
Relative LLM judge: The relative judge compares two or more LLM responses, and ranks them or selects the best one. It is commonly used in ranking-based assessments, preference modeling, and pairwise comparisons. Formally, let $(x,y_{1},y_{2})$ be a prompt-responses tuple from two judged LLMs. The relative judge maps $(x,y_{1},y_{2})$ to $(e,b)$, where $e$ is the textual evaluation of $y_{1}$ and $y_{2}$ given $x$, and $b\in\left\{0,1\right\}$ is a relative preference score. When $b=1$, the judge prefers the first response $y_{1}$; otherwise it prefers the second response $y_{2}$. We consider pairwise comparisons to simplify exposition and discuss an extension to multiple responses in Section 6.
The relative judge mitigates score inconsistencies in absolute evaluation by leveraging direct comparisons but may introduce biases due to contextual dependencies between the responses [17].
@@ 4 Quantitative LLM Judges
This work is motivated by the observation that the scores of pre-trained LLM judges (Section 3) are not calibrated to any given domain, simply because they are trained on general world knowledge. To obtain better scores, we learn to predict them from evaluations of existing judges and human scores in the domain. Our predictors are generalized linear models (GLMs) [47], which generalize linear models to non-linear functions while inheriting their efficiency. We obtain a more quantitative judge by using the predicted score from the learned model. Therefore, we call our judges quantitative.
We introduce four quantitative judges and each has the following high-level structure. The existing judge is called a base judge, and we assume that its evaluation comprises both text and score. We denote its textual evaluation by $e$ and its vector embedding by $\phi(e)\in\mathbb{R}^{d}$, where $d$ is the size of the embedding. We denote the base judge’s score by $b$. When the base judge assigns probabilities to its scores, we denote them by $p$. At inference time, when we judge, a human score is predicted from $\phi(e)$, along with $b$ or $p$. At training time, we use ground-truth human scores $s$ to train the predictor. We visualize the architectures of our quantitative judges in Figure 1 and present them next.
@@ 4.1 Least-Squares Judge
The least-squares (LS) judge is an absolute judge that predicts the score of a single response from its evaluation by a base judge. The predicted score is
@@
where $\phi(e)\in\mathbb{R}^{d}$ is the embedding of base judge’s textual evaluation, $b\in\mathbb{R}$ is its score, $u\oplus v$ is a concatenated vector of $u$ and $v$, $\theta\in\mathbb{R}^{d+1}$ is a learned model parameter, and $c$ is a population bias. The population bias plays the same role as the bias term in linear regression. We introduce $b$ so that we can always learn a judge that performs at least as well as the base judge. In particular, when $\theta=\mathbf{0}_{d}\oplus 1$ and $c=0$, (1) becomes the base judge’s score $b$.
We learn the judge by minimizing the squared loss, which lends the judge its name. Specifically, we minimize a regularized squared loss $\mathcal{L}(\theta)=\sum_{t=1}^{n}(f(e_{t},b_{t};\theta)-s_{t})^{2}+\gamma\|\theta\|_{2}^{2}$ over $n$ data points, where $e_{t}$ is the textual evaluation, $b_{t}$ is the base judge’s score, and $s_{t}$ is the human score in data point $t\in[n]$. The regularization strength $\gamma>0$ is set automatically by cross-validation.

### v2

@@ Abstract
LLM-as-a-judge is a framework where a large language model (LLM) evaluates the output of another LLM. While LLMs excel at producing qualitative textual evaluations, they often struggle to predict human preferences and numeric scores. We propose quantitative LLM judges, which align evaluation scores of existing LLM judges to humans in a given domain using regression models. The models are trained to improve the score of the original judge using its rationale and score. We present four quantitative judges for different types of absolute and relative feedback, which showcases the generality and versatility of our framework. Our framework is more computationally efficient than supervised fine-tuning and can be more statistically efficient when human feedback is limited, which is expected in practice. We validate these claims empirically on four datasets using two base judges. Our experiments show that quantitative judges can improve the predictive power of existing judges through post-hoc modeling.
@@ 1 Introduction
Measuring the quality of generated natural language presents several challenges due to the diverse range of generation methods and evaluation criteria (Gehrmann et al., 2023). The LLM-as-a-judge paradigm has recently emerged as a compelling approach to these evaluation challenges. By leveraging the reasoning capabilities of large language models (LLMs), this approach can provide more nuanced assessments that correlate better with human judgments across diverse tasks (Gu et al., 2025). LLM judges typically output both rationales and numeric scores, thus combining the comprehensiveness of human evaluation with the scalability of automated metrics.
@@
Recent studies highlight issues with LLM judges such as low alignment with human scores, miscalibration, score compression, high variance, prompt sensitivity, and leniency bias (Thakur et al., 2025; Wei et al., 2025). To address these issues, various approaches have been proposed to fine-tune LLMs to improve score prediction on specific tasks (Chiang et al., 2025; Lukasik et al., 2025). However, fine-tuning is often impractical, due to requiring large datasets and having a high computational cost. As a result, many users continue to rely on off-the-shelf models like GPT-4 or Claude (Bavaresco et al., 2024; Huang et al., 2024).
@@
The main issue with current LLM judges is that they conflate qualitative reasoning with quantitative assessment. For example, fine-tuned LLM judges like Prometheus (Kim et al., 2024b) are trained with the cross-entropy (CE) loss on gold-standard textual evaluations to generate both qualitative rationales and numeric scores. This is an inherent mismatch in objectives, as LLMs optimized for next-token prediction are tasked with producing accurate numeric scores, a fundamentally different statistical problem. While LLMs excel at producing structured textual evaluations, they are poorly suited for regressing human scores or preferences (Chiang et al., 2025).
@@
This mismatch raises a natural question: can qualitative rationales be decoupled from quantitative scores to achieve more accurate judgment? This decoupling allows the LLM produce high-quality rationales while numeric scores are predicted by classic machine learning models that are robust and loved by practitioners. This perspective is supported by prior works in interpretability and probing, which show that when model representations encode information relevant for downstream tasks, simple linear decoders can recover this information effectively (Alain and Bengio, 2017; Hewitt and Manning, 2019; Hupkes et al., 2018; Belinkov, 2022).
@@
Building on this insight, we present quantitative judges, a framework that enhances the original base judge to predict more accurate numeric scores. We introduce four different quantitative judges for absolute rating and relative preference prediction. Each judge has two stages: in the qualitative stage, a frozen LLM judge generates a rationale and initial score, and in the quantitative stage, these outputs are used to predict a better score. Our design is general, efficient, and applies to any base LLM judge. Specifically:
@@
    General: Our judges predict human scores using the base judge’s rationale and score. The predictor is a generalized linear model (GLM) (Wolke and Schwetlick, 1988) trained on human scores in the domain. The judges can be applied to absolute rating prediction, such as regression and classification, or relative preference prediction. They can also be applied to any base judge because we treat it as a black box. We show the versatility of our framework by proposing four quantitative judges.
@@
    Statistically efficient: Our judges are based on GLMs, which can be learned from limited data. This is expected in most applications of our work. Our judges are designed such that the base judge’s score or a distribution of it serves as a bias term in our model. This allows us, at least in principle, to always learn at least as good judges as the base judge.
@@
    Computationally efficient: Learning of classic machine learning models on the top of frozen LLM embeddings is more computationally efficient, and can also be more statistically efficient, than fine-tuning. We report an order of magnitude speedups in Section 5.4.
We comprehensively evaluate all proposed quantitative judges on both absolute rating and relative preference datasets. The quantitative judges consistently outperform their base judges. They can also outperform fine-tuning of the base judges on both quality and computational efficiency at the same time; and outperform an order of magnitude larger judge and specialized judges on quality. These results are complemented with additional studies on key components of our design. We conclude that quantitative judges are a practical and effective approach for improving LLM-based evaluation.
@@ 2 Background
LLMs are increasingly used not only to generate outputs but also to evaluate them, a paradigm known as LLM-as-a-judge (Kim et al., 2024b; Trivedi et al., 2024; Gu et al., 2025; Zhang et al., 2024; Thakur et al., 2025; Zhu et al., 2025). This approach uses the natural language capabilities of LLMs to simulate human evaluations, offering a scalable, cost-effective, and reproducible alternative. The output of the LLM judge can be a natural language assessment, a quantitative score, or a pairwise preference. Both the text assessment and a numeric score are often generated (Kim et al., 2024b). While efficient, scores that the LLM judges produce have been criticized as not consistently aligned with human assessments (Chiang and Lee, 2023; Gehrmann et al., 2023).
@@
LLM-as-a-judge and fine-tuning: Early methods, such as Prometheus (Kim et al., 2024b), introduced LLMs fine-tuned for both absolute rating and pairwise ranking. These models tried to replicate human judgment, often outperforming heuristic metrics in aligning with human preferences. Recent works, like LLMEval (Zhang et al., 2024) and others (Wang et al., 2024b; Li et al., 2024), further refined this setup by training on diverse instruction-response datasets, such as Alpaca52k (Bommasani et al., 2021), LMSYS-Chat (Zheng et al., 2024), and ToxicChat (Lin et al., 2023), using feedback from GPT-4 or GPT-3.5. To improve the consistency and depth of judging, CritiqueLLM (Ke et al., 2024) proposed multi-path prompting that combines pointwise and pairwise strategies and JudgeLM (Zhu et al., 2025) explored reference-free fine-tuning.
@@
Self-curation methods have been proposed to improve the alignment of judges. Yuan et al. (2024) introduced self-rewarding LMs that generate both responses and reward signals to create preference datasets for direct preference optimization (DPO) (Rafailov et al., 2023). Pang et al. (2024) and Trivedi et al. (2024) explored rationalization-based preferences by generating multiple CoT explanations from a fixed seed model. Other works (Xu et al., 2023; Zhang et al., 2023) integrated iterative feedback from stronger models or humans to refine evaluation granularity. While these methods improve alignment, they often rely on aggressive fine-tuning, which can introduce variance and bias. Small changes in the training data can lead to noticeable shifts in outputs, particularly when human feedback is limited. Moreover, strong priors from pre-training may be inadvertently distorted, undermining the consistency of judgments.
@@
LLM benchmarks and calibration challenges: Evaluating the effectiveness of LLM judges has become an active area of study. JudgeBench (Tan et al., 2024) and Eval4NLP (Zeng et al., 2024) proposed benchmarks where LLMs evaluate paired responses, aiming to reflect model reasoning and preference ranking. BigGenBench (Kim et al., 2024a) expanded this by covering $77$ diverse tasks, although with limited instances per task, affecting statistical robustness. Thakur et al. (2025) showed that while large judges align well with humans, smaller models often under-perform. To reduce evaluation variance, some methods compute mean scores by averaging outputs across multiple runs or implement voting schemes. FLEUR (Lee et al., 2024) integrated logit-based probability weighting to enhance score reliability for multi-modal evaluation. However, evaluations still suffer from dataset limitations and inconsistent generalization, especially since judges are often trained on general world knowledge and not calibrated to the particular domain being evaluated.
@@
Drawbacks of LLM-as-a-judge: Recent studies have surfaced critical limitations in LLM-as-a-judge systems. Goel et al. (2025) found that the judges tend to favor models similar in architecture or training data, creating blind spots due to shared inductive biases. Li et al. (2025a) revealed the preference leakage problem, where evaluation outcomes are skewed by overlap between judge training data and model-generated responses. The issues such as overconfidence (Tian et al., 2023; Chen et al., 2024; Xiong et al., 2024) and bias (OpenAI, 2023) persist even after extensive fine-tuning, which additionally decreases the model’s generalization (Jeong et al., 2024). Despite ongoing efforts using reinforcement learning from human feedback (RLHF), robust, unbiased, and well-calibrated evaluation remains elusive (Leng et al., 2025; Kadavath et al., 2022; Li et al., 2025b). As noted in a recent survey (Gu et al., 2025), LLM evaluators are still unreliable in high-stakes or open-ended judgment scenarios.
@@ 3 Setting
We study two types of LLM judges: for evaluating a single response and comparing two responses.
Absolute LLM judge: The absolute judge evaluates a single LLM response. The evaluation can have various forms: text, score, or both. The score can evaluate various aspects of the response, such as coherence, correctness, factual consistency, relevance, and adherence to task-specific guidelines. In this work, we assume that the judge generates both a rationale and its score. Specifically, let $(x,y)$ be a prompt-response pair from a judged LLM. The absolute judge maps $(x,y)$ to $(e,b)$, where $e$ is a rationale that judges $y$ given $x$ and $b\in\mathbb{R}$ is the associated absolute score.
The primary advantage of an absolute judge is its consistency and standardization in scoring across different responses. However, it may require extensive prompt engineering or fine-tuning to align with human scores (Kadavath et al., 2022). The relative judge, which we introduce next, mitigates it by making a direct comparison. However, it may introduce biases such as sensitivity to the order of the compared responses (Jeong et al., 2024).
Relative LLM judge: The relative judge compares two or more LLM responses, and ranks them or selects the best one. It is commonly used in ranking-based assessments, preference modeling, and pairwise comparisons. Formally, let $(x,y_{1},y_{2})$ be a prompt-responses tuple from two judged LLMs. The relative judge maps $(x,y_{1},y_{2})$ to $(e,b)$, where $e$ is the rationale that judges $y_{1}$ and $y_{2}$ given $x$, and $b\in\left\{0,1\right\}$ is the associated relative preference score. When $b=1$, the judge prefers the first response $y_{1}$; otherwise it prefers the second response $y_{2}$. We consider pairwise comparisons to simplify exposition and discuss an extension to multiple responses in Section D.1.
@@ 4 Quantitative LLM Judges
This work is motivated by the observation that the scores of pre-trained LLM judges (Section 3) are not calibrated to any given domain, simply because they are trained on general world knowledge. To obtain better scores, we learn to predict them from rationales of existing judges and human scores in the domain. Our predictors are generalized linear models (GLMs) (Wolke and Schwetlick, 1988), which generalize linear models to non-linear functions while inheriting their efficiency. We obtain a more quantitative judge by using the predicted score, and thus call our judges quantitative.
We introduce four quantitative judges and each has the following high-level structure. The existing judge is called a base judge, and we assume that its evaluation comprises both a rationale and score. We denote its rationale by $e$ and its vector embedding by $\phi(e)\in\mathbb{R}^{d}$, where $d$ is the embedding size. We denote the base judge’s score by $b$. When the base judge assigns probabilities to its scores, we denote them by $p$. At inference time, when we judge, a human score is predicted from $\phi(e)$, along with $b$ or $p$. At training time, we use ground-truth human scores to train the predictor and denote them by $s$. We show architectures of our quantitative judges in Figure 1 and present them next.
Absolute judgeRelative judgeAbsolute judgeAbsolute judgePrompt xxResponse yyPrompt xxResponse $y_{1}$Response $y_{2}$Prompt xxResponse $y_{1}$Response $y_{2}$LS / MN judgeBTL judgeBTL2 judgeText eeScore bbText eeScore bbText $e_{1}$Score $b_{1}$Text $e_{2}$Score $b_{2}$Absolute scoreRelative scoreRelative score
Figure 1: Architectures of the least-squares (LS), multinomial (MN), Bradley-Terry-Luce (BTL), and two-headed BTL (BTL2) judges described in Section 4.
@@ 4.1 Least-Squares Judge
The least-squares (LS) judge is an absolute judge that predicts the score of a single response as
@@
where $\phi(e)\in\mathbb{R}^{d}$ is the embedding of base judge’s rationale, $b\in\mathbb{R}$ is its score, $u\oplus v$ represents a concatenated vector of $u$ and $v$, $\theta\in\mathbb{R}^{d+1}$ is a learned model parameter, and $c$ is a population bias. The population bias plays the same role as the bias term in linear regression. We introduce $b$ so that we can always learn a judge that performs at least as well as the base judge. More specifically, when $\theta=\mathbf{0}_{d}\oplus 1$ and $c=0$, (1) becomes the base judge’s score $b$. While this is certainly true, we also prove in Section D.2 that this happens with a high probability when learning $\theta$.
We learn the judge by minimizing the squared loss, which lends the judge its name. Specifically, we minimize a regularized squared loss $\mathcal{L}(\theta)=\sum_{t=1}^{n}(f(e_{t},b_{t};\theta)-s_{t})^{2}+\gamma\|\theta\|_{2}^{2}$ over $n$ data points, where $e_{t}$ is the base judge’s rationale, $b_{t}$ is its score, and $s_{t}$ is the human score in data point $t\in[n]$. The regularization strength $\gamma>0$ is set automatically by cross-validation.